# Mathematical Language Models: A Survey

### v3

@@ 7. Conclusions
  • Taylor et al. (2022) Ross Taylor, Marcin Kardas, Guillem Cucurull, Thomas Scialom, Anthony Hartshorn, Elvis Saravia, Andrew Poulton, Viktor Kerkez, and Robert Stojnic. 2022. Galactica: A Large Language Model for Science. http://arxiv.org/abs/2211.09085 arXiv:2211.09085 [cs, stat].
@@
  • Thoppilan et al. (2022) Romal Thoppilan, Daniel De Freitas, Jamie Hall, Noam Shazeer, Apoorv Kulshreshtha, Heng-Tze Cheng, Alicia Jin, Taylor Bos, Leslie Baker, Yu Du, et al. 2022. Lamda: Language models for dialog applications. arXiv preprint arXiv:2201.08239 (2022).
@@
  • Touvron et al. (2023a) Hugo Touvron, Thibaut Lavril, Gautier Izacard, Xavier Martinet, Marie-Anne Lachaux, Timothée Lacroix, Baptiste Rozière, Naman Goyal, Eric Hambro, Faisal Azhar, et al. 2023a. Llama: Open and efficient foundation language models. arXiv preprint arXiv:2302.13971 (2023).
@@
  • Touvron et al. (2023b) Hugo Touvron, Louis Martin, Kevin Stone, Peter Albert, Amjad Almahairi, Yasmine Babaei, Nikolay Bashlykov, Soumya Batra, Prajjwal Bhargava, Shruti Bhosale, et al. 2023b. Llama 2: Open foundation and fine-tuned chat models. arXiv preprint arXiv:2307.09288 (2023).
@@
  • Upadhyay and Chang (2017) Shyam Upadhyay and Ming-Wei Chang. 2017. Annotating Derivations: A New Evaluation Strategy and Dataset for Algebra Word Problems. In Proceedings of the 15th Conference of the European Chapter of the Association for Computational Linguistics: Volume 1, Long Papers. 494–504.
  • Upadhyay et al. (2023) Shriyash Upadhyay, Etan Ginsberg, and Chris Callison-Burch. 2023. Improving Mathematics Tutoring With A Code Scratchpad. In Proceedings of the 18th Workshop on Innovative Use of NLP for Building Educational Applications (BEA 2023), Ekaterina Kochmar, Jill Burstein, Andrea Horbach, Ronja Laarmann-Quante, Nitin Madnani, Anaïs Tack, Victoria Yaneva, Zheng Yuan, and Torsten Zesch (Eds.). Association for Computational Linguistics, Toronto, Canada, 20–28. https://doi.org/10.18653/v1/2023.bea-1.2
  • Vaswani et al. (2017) Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones, Aidan N Gomez, Łukasz Kaiser, and Illia Polosukhin. 2017. Attention is all you need. Advances in neural information processing systems 30 (2017).
  • Wallace et al. (2019) Eric Wallace, Yizhong Wang, Sujian Li, Sameer Singh, and Matt Gardner. 2019. Do NLP Models Know Numbers? Probing Numeracy in Embeddings. In Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing and the 9th International Joint Conference on Natural Language Processing (EMNLP-IJCNLP). 5307–5315.
@@
  • Wang et al. (2021) Cunxiang Wang, Boyuan Zheng, Yuchen Niu, and Yue Zhang. 2021. Exploring Generalization Ability of Pretrained Language Models on Arithmetic and Logical Reasoning. In CCF International Conference on Natural Language Processing and Chinese Computing. 758–769.
@@
  • Wang et al. (2023a) Peiyi Wang, Lei Li, Liang Chen, Feifan Song, Binghuai Lin, Yunbo Cao, Tianyu Liu, and Zhifang Sui. 2023a. Making large language models better reasoners with alignment. arXiv preprint arXiv:2309.02144 (2023).
@@
  • Wang et al. (2023b) Xuezhi Wang, Jason Wei, Dale Schuurmans, Quoc V Le, Ed H Chi, Sharan Narang, Aakanksha Chowdhery, and Denny Zhou. 2023b. Self-Consistency Improves Chain of Thought Reasoning in Language Models. In The Eleventh International Conference on Learning Representations.
@@
  • Wang et al. (2017) Yan Wang, Xiaojiang Liu, and Shuming Shi. 2017. Deep neural solver for math word problems. In Proceedings of the 2017 conference on empirical methods in natural language processing. 845–854.
@@
  • Wang et al. (2023c) Zengzhi Wang, Rui Xia, and Pengfei Liu. 2023c. Generative AI for Math: Part I–MathPile: A Billion-Token-Scale Pretraining Corpus for Math. arXiv preprint arXiv:2312.17120 (2023).
@@
  • Wei et al. (2022) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Fei Xia, Ed Chi, Quoc V Le, Denny Zhou, et al. 2022. Chain-of-thought prompting elicits reasoning in large language models. Advances in Neural Information Processing Systems 35 (2022), 24824–24837.
@@
  • Welleck et al. (2021) Sean Welleck, Jiacheng Liu, Ronan Le Bras, Hannaneh Hajishirzi, Yejin Choi, and Kyunghyun Cho. 2021. NaturalProofs: Mathematical Theorem Proving in Natural Language. In Thirty-fifth Conference on Neural Information Processing Systems Datasets and Benchmarks Track (Round 1).
@@
  • Welleck et al. (2022) Sean Welleck, Jiacheng Liu, Ximing Lu, Hannaneh Hajishirzi, and Yejin Choi. 2022. Naturalprover: Grounded mathematical proof generation with language models. Advances in Neural Information Processing Systems 35 (2022), 4913–4927.
@@
  • Welleck and Saha (2023) Sean Welleck and Rahul Saha. 2023. llmstep: LLM proofstep suggestions in Lean. https://github.com/wellecks/llmstep.
@@
  • Wenzel et al. (2008) Makarius Wenzel, Lawrence C Paulson, and Tobias Nipkow. 2008. The isabelle framework. In Theorem Proving in Higher Order Logics: 21st International Conference, TPHOLs 2008, Montreal, Canada, August 18-21, 2008. Proceedings 21. Springer, 33–38.
@@
  • Wu et al. (2021) Yuhuai Wu, Albert Jiang, Jimmy Ba, and Roger Baker Grosse. 2021. INT: An Inequality Benchmark for Evaluating Generalization in Theorem Proving. In International Conference on Learning Representations.
@@
  • Wu et al. (2022) Yuhuai Wu, Albert Qiaochu Jiang, Wenda Li, Markus N. Rabe, Charles Staats, Mateja Jamnik, and Christian Szegedy. 2022. Autoformalization with Large Language Models. In NeurIPS. http://papers.nips.cc/paper_files/paper/2022/hash/d0c6bc641a56bebee9d985b937307367-Abstract-Conference.html
@@
  • Xu et al. (2023) Can Xu, Qingfeng Sun, Kai Zheng, Xiubo Geng, Pu Zhao, Jiazhan Feng, Chongyang Tao, and Daxin Jiang. 2023. WizardLM: Empowering Large Language Models to Follow Complex Instructions. http://arxiv.org/abs/2304.12244 arXiv:2304.12244 [cs].
@@
  • Yang and Deng (2019) Kaiyu Yang and Jia Deng. 2019. Learning to prove theorems via interacting with proof assistants. In International Conference on Machine Learning. PMLR, 6984–6994.
@@
  • Yang et al. (2023) Zhen Yang, Ming Ding, Qingsong Lv, Zhihuan Jiang, Zehai He, Yuyi Guo, Jinfeng Bai, and Jie Tang. 2023. GPT Can Solve Mathematical Problems Without a Calculator. arXiv preprint arXiv:2309.03241 (2023).
@@
  • Yao et al. (2023b) Shunyu Yao, Dian Yu, Jeffrey Zhao, Izhak Shafran, Thomas L Griffiths, Yuan Cao, and Karthik Narasimhan. 2023b. Tree of thoughts: Deliberate problem solving with large language models. arXiv preprint arXiv:2305.10601 (2023).
@@
  • Yao et al. (2023a) Yao Yao, Zuchao Li, and Hai Zhao. 2023a. Beyond Chain-of-Thought, Effective Graph-of-Thought Reasoning in Large Language Models. arXiv preprint arXiv:2305.16582 (2023).
@@
  • Yoran et al. (2023) Ori Yoran, Tomer Wolfson, Ben Bogin, Uri Katz, Daniel Deutch, and Jonathan Berant. 2023. Answering questions by meta-reasoning over multiple chains of thought. arXiv preprint arXiv:2304.13007 (2023).
@@
  • Yu et al. (2023a) Junchi Yu, Ran He, and Rex Ying. 2023a. Thought Propagation: An Analogical Approach to Complex Reasoning with Large Language Models. arXiv preprint arXiv:2310.03965 (2023).
@@
  • Yu et al. (2023b) Longhui Yu, Weisen Jiang, Han Shi, Jincheng Yu, Zhengying Liu, Yu Zhang, James T Kwok, Zhenguo Li, Adrian Weller, and Weiyang Liu. 2023b. Metamath: Bootstrap your own mathematical questions for large language models. arXiv preprint arXiv:2309.12284 (2023).
@@
  • Yuan et al. (2023a) Zheng Yuan, Hongyi Yuan, Chengpeng Li, Guanting Dong, Chuanqi Tan, and Chang Zhou. 2023a. Scaling relationship on learning mathematical reasoning with large language models. arXiv preprint arXiv:2308.01825 (2023).
@@
  • Yuan et al. (2023b) Zheng Yuan, Hongyi Yuan, Chuanqi Tan, Wei Wang, and Songfang Huang. 2023b. How well do Large Language Models perform in Arithmetic tasks? arXiv preprint arXiv:2304.02015 (2023).
@@
  • Yue et al. (2023) Xiang Yue, Xingwei Qu, Ge Zhang, Yao Fu, Wenhao Huang, Huan Sun, Yu Su, and Wenhu Chen. 2023. Mammoth: Building math generalist models through hybrid instruction tuning. arXiv preprint arXiv:2309.05653 (2023).
@@
  • Zeng et al. (2022) Andy Zeng, Maria Attarian, Brian Ichter, Krzysztof Choromanski, Adrian Wong, Stefan Welker, Federico Tombari, Aveek Purohit, Michael Ryoo, Vikas Sindhwani, et al. 2022. Socratic models: Composing zero-shot multimodal reasoning with language. arXiv preprint arXiv:2204.00598 (2022).
@@
  • Zhang et al. (2022) Ming-Liang Zhang, Fei Yin, Yi-Han Hao, and Cheng-Lin Liu. 2022. Plane Geometry Diagram Parsing. In Proceedings of the Thirty-First International Joint Conference on Artificial Intelligence, IJCAI-22. 1636–1643. https://doi.org/10.24963/ijcai.2022/228
@@
  • Zhang et al. (2023b) Ming-Liang Zhang, Fei Yin, and Cheng-Lin Liu. 2023b. A Multi-Modal Neural Geometric Solver with Textual Clauses Parsed from Diagram. In IJCAI.
@@
  • Zhang et al. (2023a) Shengyu Zhang, Linfeng Dong, Xiaoya Li, Sen Zhang, Xiaofei Sun, Shuhe Wang, Jiwei Li, Runyi Hu, Tianwei Zhang, Fei Wu, and Guoyin Wang. 2023a. Instruction Tuning for Large Language Models: A Survey. http://arxiv.org/abs/2308.10792 arXiv:2308.10792 [cs].
@@
  • Zhang et al. (2020a) Xikun Zhang, Deepak Ramachandran, Ian Tenney, Yanai Elazar, and Dan Roth. 2020a. Do language embeddings capture scales?. In Findings of the Association for Computational Linguistics: EMNLP 2020, Trevor Cohn, Yulan He, and Yang Liu (Eds.). Association for Computational Linguistics, 4889–4896. https://doi.org/10.18653/v1/2020.findings-emnlp.439
@@
  • Zhang et al. (2020b) Xikun Zhang, Deepak Ramachandran, Ian Tenney, Yanai Elazar, and Dan Roth. 2020b. Do Language Embeddings capture Scales?. In Proceedings of the Third BlackboxNLP Workshop on Analyzing and Interpreting Neural Networks for NLP. 292–299.
@@
  • Zhang et al. (2023c) Zhuosheng Zhang, Aston Zhang, Mu Li, and Alex Smola. 2023c. Automatic Chain of Thought Prompting in Large Language Models. In The Eleventh International Conference on Learning Representations.
@@
  • Zhao et al. (2023a) Ruochen Zhao, Xingxuan Li, Shafiq Joty, Chengwei Qin, and Lidong Bing. 2023a. Verify-and-edit: A knowledge-enhanced chain-of-thought framework. arXiv preprint arXiv:2305.03268 (2023).
  • Zhao et al. (2020) Wei Zhao, Mingyue Shang, Yang Liu, Liang Wang, and Jingming Liu. 2020. Ape210k: A large-scale and template-rich dataset of math word problems. arXiv preprint arXiv:2009.11506 (2020).
@@
  • Zhao et al. (2023b) Wayne Xin Zhao, Kun Zhou, Junyi Li, Tianyi Tang, Xiaolei Wang, Yupeng Hou, Yingqian Min, Beichen Zhang, Junjie Zhang, Zican Dong, et al. 2023b. A survey of large language models. arXiv preprint arXiv:2303.18223 (2023).
@@
  • Zhao et al. (2022) Yilun Zhao, Yunxiang Li, Chenying Li, and Rui Zhang. 2022. MultiHiertt: Numerical Reasoning over Multi Hierarchical Tabular and Textual Data. In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers). 6588–6600.
@@
  • Zheng et al. (2023) Kunhao Zheng, Jesse Michael Han, and Stanislas Polu. 2023. miniF2F: a cross-system benchmark for formal Olympiad-level mathematics. In International Conference on Learning Representations.

### v4

@@ Acknowledgements.
  • Taylor et al. (2022) Ross Taylor, Marcin Kardas, Guillem Cucurull, Thomas Scialom, Anthony Hartshorn, Elvis Saravia, Andrew Poulton, Viktor Kerkez, and Robert Stojnic. 2022. Galactica: A Large Language Model for Science.
  • Team et al. (2023) Gemini Team, Rohan Anil, Sebastian Borgeaud, Jean-Baptiste Alayrac, Jiahui Yu, Radu Soricut, Johan Schalkwyk, Andrew M Dai, Anja Hauth, Katie Millican, et al. 2023. Gemini: a family of highly capable multimodal models. arXiv (2023).
  • Team (2024) Qwen Team. 2024. QwQ: Reflect Deeply on the Boundaries of the Unknown. https://qwenlm.github.io/blog/qwq-32b-preview/
@@
  • Thoppilan et al. (2022) Romal Thoppilan, Daniel De Freitas, Jamie Hall, Noam Shazeer, Apoorv Kulshreshtha, Heng-Tze Cheng, Alicia Jin, Taylor Bos, Leslie Baker, Yu Du, et al. 2022. Lamda: Language models for dialog applications. arXiv (2022).
@@
  • Touvron et al. (2023a) Hugo Touvron, Thibaut Lavril, Gautier Izacard, Xavier Martinet, Marie-Anne Lachaux, Timothée Lacroix, Baptiste Rozière, Naman Goyal, Eric Hambro, Faisal Azhar, et al. 2023a. Llama: Open and efficient foundation language models. arXiv (2023).
@@
  • Touvron et al. (2023b) Hugo Touvron, Louis Martin, Kevin Stone, Peter Albert, Amjad Almahairi, Yasmine Babaei, Nikolay Bashlykov, Soumya Batra, Prajjwal Bhargava, Shruti Bhosale, et al. 2023b. Llama 2: Open foundation and fine-tuned chat models. arXiv (2023).
  • Trung et al. (2024) Luong Trung, Xinbo Zhang, Zhanming Jie, Peng Sun, Xiaoran Jin, and Hang Li. 2024. Reft: Reasoning with reinforced fine-tuning. In ACL. 7601–7614.
@@
  • Upadhyay and Chang (2017) Shyam Upadhyay and Ming-Wei Chang. 2017. Annotating Derivations: A New Evaluation Strategy and Dataset for Algebra Word Problems. In EACL. 494–504.
  • Upadhyay et al. (2023) Shriyash Upadhyay, Etan Ginsberg, and Chris Callison-Burch. 2023. Improving Mathematics Tutoring With A Code Scratchpad. In BEA. 20–28.
  • Vaswani et al. (2017) Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones, Aidan N Gomez, Łukasz Kaiser, and Illia Polosukhin. 2017. Attention is all you need. NeurIPS 30 (2017).
  • Wallace et al. (2019) Eric Wallace, Yizhong Wang, Sujian Li, Sameer Singh, and Matt Gardner. 2019. Do NLP Models Know Numbers? Probing Numeracy in Embeddings. In EMNLP. 5307–5315.
@@
  • Wang et al. (2021) Cunxiang Wang, Boyuan Zheng, Yuchen Niu, and Yue Zhang. 2021. Exploring Generalization Ability of Pretrained Language Models on Arithmetic and Logical Reasoning. In NLPCC. 758–769.
  • Wang et al. (2024b) Ke Wang, Junting Pan, Weikang Shi, Zimu Lu, Mingjie Zhan, and Hongsheng Li. 2024b. Measuring multimodal mathematical reasoning with math-vision dataset. arXiv (2024).
  • Wang et al. (2024a) Peng Wang, Shuai Bai, Sinan Tan, Shijie Wang, Zhihao Fan, Jinze Bai, Keqin Chen, Xuejing Liu, Jialin Wang, Wenbin Ge, Yang Fan, Kai Dang, Mengfei Du, Xuancheng Ren, Rui Men, Dayiheng Liu, Chang Zhou, Jingren Zhou, and Junyang Lin. 2024a. Qwen2-VL: Enhancing Vision-Language Model’s Perception of the World at Any Resolution. arXiv (2024).
@@
  • Wang et al. (2023a) Peiyi Wang, Lei Li, Liang Chen, Feifan Song, Binghuai Lin, Yunbo Cao, Tianyu Liu, and Zhifang Sui. 2023a. Making large language models better reasoners with alignment. arXiv (2023).
  • Wang et al. (2023b) Weihan Wang, Qingsong Lv, Wenmeng Yu, Wenyi Hong, Ji Qi, Yan Wang, Junhui Ji, Zhuoyi Yang, Lei Zhao, Xixuan Song, Jiazheng Xu, Bin Xu, Juanzi Li, Yuxiao Dong, Ming Ding, and Jie Tang. 2023b. CogVLM: Visual Expert for Pretrained Language Models. arXiv:2311.03079
  • Wang et al. (2023c) Xingyao Wang, Zihan Wang, Jiateng Liu, Yangyi Chen, Lifan Yuan, Hao Peng, and Heng Ji. 2023c. Mint: Evaluating llms in multi-turn interaction with tools and language feedback. arXiv (2023).
@@
  • Wang et al. (2023d) Xuezhi Wang, Jason Wei, Dale Schuurmans, Quoc V Le, Ed H Chi, Sharan Narang, Aakanksha Chowdhery, and Denny Zhou. 2023d. Self-Consistency Improves Chain of Thought Reasoning in Language Models. In ICLR.
  • wang and Fu (2024) Yifan wang and Yun Fu. 2024. Understanding, Abstracting and Checking: Evoking Complicated Multimodal Reasoning in LMMs.
@@
  • Wang et al. (2017) Yan Wang, Xiaojiang Liu, and Shuming Shi. 2017. Deep neural solver for math word problems. In EMNLP. 845–854.
@@
  • Wang et al. (2023e) Zengzhi Wang, Rui Xia, and Pengfei Liu. 2023e. Generative AI for Math: Part I–MathPile: A Billion-Token-Scale Pretraining Corpus for Math. arXiv (2023).
@@
  • Wei et al. (2022) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Fei Xia, Ed Chi, Quoc V Le, Denny Zhou, et al. 2022. Chain-of-thought prompting elicits reasoning in large language models. NeurIPS 35 (2022), 24824–24837.
@@
  • Welleck et al. (2021) Sean Welleck, Jiacheng Liu, Ronan Le Bras, Hannaneh Hajishirzi, Yejin Choi, and Kyunghyun Cho. 2021. NaturalProofs: Mathematical Theorem Proving in Natural Language. In NeurIPS Datasets and Benchmarks Track (Round 1).
@@
  • Welleck et al. (2022) Sean Welleck, Jiacheng Liu, Ximing Lu, Hannaneh Hajishirzi, and Yejin Choi. 2022. Naturalprover: Grounded mathematical proof generation with language models. NeurIPS 35 (2022), 4913–4927.
@@
  • Welleck and Saha (2023) Sean Welleck and Rahul Saha. 2023. llmstep: LLM proofstep suggestions in Lean. In NeurIPS.
@@
  • Wenzel et al. (2008) Makarius Wenzel, Lawrence C Paulson, and Tobias Nipkow. 2008. The isabelle framework. In TPHOLs. 33–38.
@@
  • Wu et al. (2021) Yuhuai Wu, Albert Jiang, Jimmy Ba, and Roger Baker Grosse. 2021. INT: An Inequality Benchmark for Evaluating Generalization in Theorem Proving. In ICLR.
@@
  • Wu et al. (2022) Yuhuai Wu, Albert Qiaochu Jiang, Wenda Li, Markus N. Rabe, Charles Staats, Mateja Jamnik, and Christian Szegedy. 2022. Autoformalization with Large Language Models. In NeurIPS.
@@
  • Xia et al. (2024) Shijie Xia, Xuefeng Li, Yixin Liu, Tongshuang Wu, and Pengfei Liu. 2024. Evaluating Mathematical Reasoning Beyond Accuracy. arXiv (2024).
  • Xiang et al. (2024) Kun Xiang, Zhili Liu, Zihao Jiang, Yunshuang Nie, Runhui Huang, Haoxiang Fan, Hanhui Li, Weiran Huang, Yihan Zeng, Jianhua Han, et al. 2024. AtomThink: A Slow Thinking Framework for Multimodal Mathematical Reasoning. arXiv (2024).
@@
  • Xu et al. (2023) Can Xu, Qingfeng Sun, Kai Zheng, Xiubo Geng, Pu Zhao, Jiazhan Feng, Chongyang Tao, and Daxin Jiang. 2023. Wizardlm: Empowering large language models to follow complex instructions. arXiv (2023).
  • Xu et al. (2024b) Guowei Xu, Peng Jin, Li Hao, Yibing Song, Lichao Sun, and Li Yuan. 2024b. LLaVA-o1: Let Vision Language Models Reason Step-by-Step. arXiv (2024).
  • Xu et al. (2024a) Jundong Xu, Hao Fei, Liangming Pan, Qian Liu, Mong-Li Lee, and Wynne Hsu. 2024a. Faithful Logical Reasoning via Symbolic Chain-of-Thought. arXiv (2024).
  • Yan et al. (2024) Yibo Yan, Jiamin Su, Jianxiang He, Fangteng Fu, Xu Zheng, Yuanhuiyi Lyu, Kun Wang, Shen Wang, Qingsong Wen, and Xuming Hu. 2024. A Survey of Mathematical Reasoning in the Era of Multimodal Large Language Model: Benchmark, Method & Challenges. arXiv (2024).
  • Yang et al. (2024) An Yang, Beichen Zhang, Binyuan Hui, Bofei Gao, Bowen Yu, Chengpeng Li, Dayiheng Liu, Jianhong Tu, Jingren Zhou, Junyang Lin, et al. 2024. Qwen2. 5-math technical report: Toward mathematical expert model via self-improvement. arXiv (2024).
@@
  • Yang and Deng (2019) Kaiyu Yang and Jia Deng. 2019. Learning to prove theorems via interacting with proof assistants. In ICML. 6984–6994.
@@
  • Yang et al. (2023) Zhen Yang, Ming Ding, Qingsong Lv, Zhihuan Jiang, Zehai He, Yuyi Guo, Jinfeng Bai, and Jie Tang. 2023. GPT Can Solve Mathematical Problems Without a Calculator. arXiv (2023).
  • Yang et al. (2022) Zhicheng Yang, Jinghui Qin, Jiaqi Chen, Liang Lin, and Xiaodan Liang. 2022. Logicsolver: Towards interpretable math word problem solving with logical prompt-enhanced learning. arXiv (2022).
@@
  • Yao et al. (2023b) Shunyu Yao, Dian Yu, Jeffrey Zhao, Izhak Shafran, Thomas L Griffiths, Yuan Cao, and Karthik Narasimhan. 2023b. Tree of thoughts: Deliberate problem solving with large language models. arXiv (2023).
@@
  • Yao et al. (2023a) Yao Yao, Zuchao Li, and Hai Zhao. 2023a. Beyond Chain-of-Thought, Effective Graph-of-Thought Reasoning in Large Language Models. arXiv (2023).
  • Yin et al. (2024) Shuo Yin, Weihao You, Zhilong Ji, Guoqiang Zhong, and Jinfeng Bai. 2024. MuMath-Code: Combining Tool-Use Large Language Models with Multi-perspective Data Augmentation for Mathematical Reasoning. arXiv (2024).
  • Ying et al. (2024) Huaiyuan Ying, Shuo Zhang, Linyang Li, Zhejian Zhou, Yunfan Shao, Zhaoye Fei, Yichuan Ma, Jiawei Hong, Kuikun Liu, Ziyi Wang, Yudong Wang, Zijian Wu, Shuaibin Li, Fengzhe Zhou, Hongwei Liu, Songyang Zhang, Wenwei Zhang, Hang Yan, Xipeng Qiu, Jiayu Wang, Kai Chen, and Dahua Lin. 2024. InternLM-Math: Open Math Large Language Models Toward Verifiable Reasoning.
@@
  • Yoran et al. (2023) Ori Yoran, Tomer Wolfson, Ben Bogin, Uri Katz, Daniel Deutch, and Jonathan Berant. 2023. Answering questions by meta-reasoning over multiple chains of thought. arXiv (2023).
@@
  • Yu et al. (2023a) Junchi Yu, Ran He, and Rex Ying. 2023a. Thought Propagation: An Analogical Approach to Complex Reasoning with Large Language Models. arXiv (2023).
@@
  • Yu et al. (2023b) Longhui Yu, Weisen Jiang, Han Shi, Jincheng Yu, Zhengying Liu, Yu Zhang, James T Kwok, Zhenguo Li, Adrian Weller, and Weiyang Liu. 2023b. Metamath: Bootstrap your own mathematical questions for large language models. arXiv (2023).
  • Yuan et al. (2024) Lifan Yuan, Ganqu Cui, Hanbin Wang, Ning Ding, Xingyao Wang, Jia Deng, Boji Shan, Huimin Chen, Ruobing Xie, Yankai Lin, et al. 2024. Advancing llm reasoning generalists with preference trees. arXiv (2024).
@@
  • Yuan et al. (2023a) Zheng Yuan, Hongyi Yuan, Chengpeng Li, Guanting Dong, Chuanqi Tan, and Chang Zhou. 2023a. Scaling relationship on learning mathematical reasoning with large language models. arXiv (2023).
@@
  • Yuan et al. (2023b) Zheng Yuan, Hongyi Yuan, Chuanqi Tan, Wei Wang, and Songfang Huang. 2023b. How well do Large Language Models perform in Arithmetic tasks? arXiv (2023).
  • Yue et al. (2024) Xiang Yue, Yuansheng Ni, Kai Zhang, Tianyu Zheng, Ruoqi Liu, Ge Zhang, Samuel Stevens, Dongfu Jiang, Weiming Ren, Yuxuan Sun, Cong Wei, Botao Yu, Ruibin Yuan, Renliang Sun, Ming Yin, Boyuan Zheng, Zhenzhu Yang, Yibo Liu, Wenhao Huang, Huan Sun, Yu Su, and Wenhu Chen. 2024. MMMU: A Massive Multi-discipline Multimodal Understanding and Reasoning Benchmark for Expert AGI. arXiv:2311.16502
@@
  • Yue et al. (2023) Xiang Yue, Xingwei Qu, Ge Zhang, Yao Fu, Wenhao Huang, Huan Sun, Yu Su, and Wenhu Chen. 2023. Mammoth: Building math generalist models through hybrid instruction tuning. arXiv (2023).
  • Zelikman et al. (2024) Eric Zelikman, Georges Harik, Yijia Shao, Varuna Jayasiri, Nick Haber, and Noah D Goodman. 2024. Quiet-star: Language models can teach themselves to think before speaking. arXiv (2024).
  • Zelikman et al. (2022) Eric Zelikman, Yuhuai Wu, Jesse Mu, and Noah Goodman. 2022. Star: Bootstrapping reasoning with reasoning. NeurIPS 35 (2022), 15476–15488.
@@
  • Zeng et al. (2022) Andy Zeng, Maria Attarian, Brian Ichter, Krzysztof Choromanski, Adrian Wong, Stefan Welker, Federico Tombari, Aveek Purohit, Michael Ryoo, Vikas Sindhwani, et al. 2022. Socratic models: Composing zero-shot multimodal reasoning with language. arXiv (2022).
  • Zhang et al. (2024b) Di Zhang, Xiaoshui Huang, Dongzhan Zhou, Yuqiang Li, and Wanli Ouyang. 2024b. Accessing gpt-4 level mathematical olympiad solutions via monte carlo tree self-refine with llama-3 8b. arXiv (2024).
  • Zhang et al. (2023b) Mengxue Zhang, Zichao Wang, Zhichao Yang, Weiqi Feng, and Andrew Lan. 2023b. Interpretable math word problem solution generation via step-by-step planning. arXiv (2023).
@@
  • Zhang et al. (2022) Ming-Liang Zhang, Fei Yin, Yi-Han Hao, and Cheng-Lin Liu. 2022. Plane Geometry Diagram Parsing. In IJCAI. 1636–1643.
@@
  • Zhang et al. (2023d) Ming-Liang Zhang, Fei Yin, and Cheng-Lin Liu. 2023d. A multi-modal neural geometric solver with textual clauses parsed from diagram. In IJCAI. 3374–3382.
  • Zhang et al. (2025) Renrui Zhang, Dongzhi Jiang, Yichi Zhang, Haokun Lin, Ziyu Guo, Pengshuo Qiu, Aojun Zhou, Pan Lu, Kai-Wei Chang, Yu Qiao, et al. 2025. Mathverse: Does your multi-modal llm truly see the diagrams in visual math problems?. In ECCV. 169–186.
@@
  • Zhang et al. (2023a) Shengyu Zhang, Linfeng Dong, Xiaoya Li, Sen Zhang, Xiaofei Sun, Shuhe Wang, Jiwei Li, Runyi Hu, Tianwei Zhang, Fei Wu, et al. 2023a. Instruction tuning for large language models: A survey. arXiv (2023).
@@
  • Zhang et al. (2020a) Xikun Zhang, Deepak Ramachandran, Ian Tenney, Yanai Elazar, and Dan Roth. 2020a. Do language embeddings capture scales?. In EMNLP. 4889–4896.
@@
  • Zhang et al. (2020b) Xikun Zhang, Deepak Ramachandran, Ian Tenney, Yanai Elazar, and Dan Roth. 2020b. Do Language Embeddings capture Scales?. In BlackboxNLP Workshop on Analyzing and Interpreting Neural Networks for NLP. 292–299.
  • Zhang et al. (2023c) Yifan Zhang, Jingqin Yang, Yang Yuan, and Andrew Chi-Chih Yao. 2023c. Cumulative reasoning with large language models. arXiv (2023).
  • Zhang et al. (2024a) Zhihan Zhang, Tao Ge, Zhenwen Liang, Wenhao Yu, Dian Yu, Mengzhao Jia, Dong Yu, and Meng Jiang. 2024a. Learn Beyond The Answer: Training Language Models with Reflection for Mathematical Reasoning. arXiv (2024).
@@
  • Zhang et al. (2023e) Zhuosheng Zhang, Aston Zhang, Mu Li, and Alex Smola. 2023e. Automatic Chain of Thought Prompting in Large Language Models. In ICLR.
  • Zhao et al. (2024b) Jun Zhao, Jingqi Tong, Yurong Mou, Ming Zhang, Qi Zhang, and Xuan-Jing Huang. 2024b. Exploring the Compositional Deficiency of Large Language Models in Mathematical Reasoning Through Trap Problems. In EMNLP. 16361–16376.
@@
  • Zhao et al. (2023a) Ruochen Zhao, Xingxuan Li, Shafiq Joty, Chengwei Qin, and Lidong Bing. 2023a. Verify-and-edit: A knowledge-enhanced chain-of-thought framework. arXiv (2023).
@@
  • Zhao et al. (2023b) Wayne Xin Zhao, Kun Zhou, Junyi Li, Tianyi Tang, Xiaolei Wang, Yupeng Hou, Yingqian Min, Beichen Zhang, Junjie Zhang, Zican Dong, et al. 2023b. A survey of large language models. arXiv (2023).
@@
  • Zhao et al. (2022) Yilun Zhao, Yunxiang Li, Chenying Li, and Rui Zhang. 2022. MultiHiertt: Numerical Reasoning over Multi Hierarchical Tabular and Textual Data. In ACL. 6588–6600.
  • Zhao et al. (2024c) Yu Zhao, Huifeng Yin, Bo Zeng, Hao Wang, Tianqi Shi, Chenyang Lyu, Longyue Wang, Weihua Luo, and Kaifu Zhang. 2024c. Marco-o1: Towards open reasoning models for open-ended solutions. arXiv (2024).
  • Zhao et al. (2024a) Zilong Zhao, Yao Rong, Dongyang Guo, Emek Gözlüklü, Emir Gülboy, and Enkelejda Kasneci. 2024a. Stepwise Self-Consistent Mathematical Reasoning with Large Language Models. arXiv (2024).
@@
  • Zheng et al. (2023) Kunhao Zheng, Jesse Michael Han, and Stanislas Polu. 2023. miniF2F: a cross-system benchmark for formal Olympiad-level mathematics. In ICLR.